OpenStudy (anonymous):

The solutions of x^3-6x^2-19x+84=0 are x=-3,4,7 Write the polynomial in factored form

myininaya (myininaya):

You can find factors by solving each x=i for 0 by subtracting i on both side like x-i=0 or if you have x=-i then x+i=0

myininaya (myininaya):

So then you could say x-i or x+i are factors

OpenStudy (anonymous):

Oh so I should write x-3 etc

myininaya (myininaya):

x=-3 means x+3=0 not x-3=0 x=-3 add 3 on both sides x+3=-3+3 x+3=0 so x+3 is a factor

OpenStudy (anonymous):

Oh my bad I meant to write x+3